OpenStudy (anonymous):

Write an equation in slope intercept form of the line satisfying the given conditions.The line passes through (3,-7) and is paralleled to the line whose equation is y=4x+2

OpenStudy (anonymous):

What's the slope of the line with the equation y=4x+2?

OpenStudy (anonymous):

slope 4, y intercept 2

OpenStudy (anonymous):

Right, and parallel lines have the same slope. So use that info to get the equation. you have slope 4 and point (3,-7)

OpenStudy (anonymous):

y=4x-7?

OpenStudy (anonymous):

Nope. use y-y1=m(x-x1)

OpenStudy (anonymous):

So where would the slope go?

OpenStudy (anonymous):

m=slope (x1,y1) is the point you have y+7=4(x-3) now simplify and isolate y